OpenStudy (anonymous):

Evaluate the integral using the given substitution. ∫dx/(sqrt 4x+4), u=4x+4 Need help with steps, will award medal!

OpenStudy (anonymous):



OpenStudy (anonymous):

ok so you are even given the substitution , find derivative of u , then substitude in the integral

OpenStudy (anonymous):

Derivative of u is 4

OpenStudy (anonymous):

wait you mean antiderivative?

OpenStudy (anonymous):

no no i mean derivative , 4 is correct but let me write it for you du=4dx

OpenStudy (anonymous):

ok so now \[dx=\frac{ du }{ 4 }\] and u=4x+4 so substitude in the integral for that

OpenStudy (anonymous):

Okay got it, then what

OpenStudy (anonymous):

so what do you get when u substitude for u and replace dx with du/4

OpenStudy (anonymous):

I need you to set that up for me, I'm a little confused

OpenStudy (anonymous):

ok so we have u=4x+4 and dx=du/4 so the integral becomes \[\int\limits_{}^{}{\frac{ du }{ 4\sqrt{u} }}\]

OpenStudy (anonymous):

do you know to solve this integral now ?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

huh , ok so the constant there we can pull out of the integral so it would be